# Long-fiber Sagnac interferometers for twin field quantum key distribution networks

### v2

@@ Abstract
A Sagnac loop structure can help overcome the major difficulty in the practical implementation of a twin field quantum key distribution (TFQKD) network, namely, the need to stabilize the phase of a quantum state over many kilometers of fiber. Unfortunately, Rayleigh backscattering noise limits the signal-to-noise ratio for Sagnac systems containing long fibers and lossy photonic devices. Here, we solve this problem by sending optical pulses in long on-off bursts and using time post-selection on measurements taken with free-run single-photon avalanche detectors. We also investigate the impact of the residual phase noise uncompensated by the Sagnac structure and find that the variance of the phase noise scales as loop length to the third power, verifying an existing calculation in the literature. We measure the interference visibility in Sagnac loops of varying length without active phase or polarization stabilization and achieve $>97\%$ visibility in 200 km ultra-low-loss fiber, which is, to our knowledge, the longest fiber Sagnac interferometer demonstrated. Our results indicate the suitability of a Sagnac system for long-distance TFQKD networks, an important step towards the practical implementation of metropolitan quantum networks.
@@ I Introduction
In TFQKD, two remote users (conventionally called Alice and Bob) send coherent pulses encoded with an optical phase to an untrusted central node (conventionally called Charlie) who performs a single-photon interference measurement on the states [16], which requires phase stability between the two interfering optical paths. In most TFQKD experiments, Alice’s and Bob’s laser sources are phase-locked to a reference laser over additional fiber channels, forming a large Mach-Zehnder interferometer [17, 18, 19, 20, 21]. Recently, there have also been demonstrations without phase locking the two sources [22, 23]. In all of the aforementioned works, however, expensive superconducting nanowire single-photon detectors with precise timing resolution are used in order to measure [17, 19, 21, 22] and/or compensate [18, 20, 23] phase noise. Such configurations are inherently impractical for a TFQKD network, where more than one user pair share Charlie’s measurement station. Moreover, unless employing sub-Hz linewidth lasers with matched frequencies as in Ref. [23], it is practically very difficult to stabilize interferometers with asymmetric fiber links, a situation that is inevitable in a multi-user-pair TFQKD network.
Unlike the other systems, TFQKD based on a Sagnac interferometer shows promise for easy networking and low-cost implementation [24, 25, 26] due to common-path stable interference. However, Sagnac TFQKD with auto phase stabilization has been restricted to fiber distances of less than 20 km. There are two main challenges to implementing Sagnac TFQKD over long fibers. One is the uncompensated fiber phase noise due to temperature fluctuations and vibrations in the surrounding environment that cause the phase in the fiber to fluctuate during the loop transit time. The other is the Rayleigh backscattering noise in fiber which is problematic due to the bidirectional transmission in a Sagnac loop. This noise is proportional to the intensity and repetition rate of the source at Charlie and thus, for a high-speed and long-distance Sagnac TFQKD network (which contains several lossy photonic devices), the signal-to-noise ratio (SNR) at the detectors becomes limited by the backscattering.
@@
Indeed, no existing literature assesses the feasibility of implementing high-speed long-distance Sagnac TFQKD, with respect to auto phase stability as well as Rayleigh backscattering noise. We overcome both the phase noise and the Rayleigh backscattering noise in fiber to achieve stable interference in loop lengths reaching 200 km. We mitigate the fiber phase noise via thermal and vibrational isolation of the fiber spools; no active phase or polarization stabilization is performed, nor do we perform phase post-compensation. We report measurements on the phase noise as a function of Sagnac loop length, $L$, and find the variance to scale roughly as $L^{3}$. We circumvent the fiber backscattering noise by sending pulses in long, on-off bursts and using post-selection on time-resolved measurements taken with free-run single-photon avalanche detectors (SPADs). We achieve $>99\%$ visibility in 150 km fiber, and $>97\%$ visibility in 200 km. Our results represent record performance in terms of Sagnac interference over long fibers, and indicate that the Sagnac loop is imminently suitable as a platform for long-distance TFQKD networks.
@@ II Sagnac Interferometry in Fiber
Figure 1: Sagnac interferometer experimental setup. Light from a continuous-wave laser launches into a fiber ring via a 50:50 beam splitter (BS). Interference of the counterpropagating paths is measured by detectors $D_{0}$ and $D_{1}$. Components on the hexagonal or square platforms are added when measuring phase noise or visibility, respectively. Classical photodetectors and single-photon avalanche detectors are used for phase noise and visibility measurements, respectively. IM: intensity modulator; VOA: variable optical attenuator; C: circulator; PC: polarization controller; PM: phase modulator; RFSA: rf spectrum analyzer; Osc: oscilloscope.
@@ IV Fiber Phase Noise
$I(\delta\phi(t))$ was measured for loop lengths ranging from five to 125 km. Spools consisted of SMF-28 fiber (fiber attenuation coefficient, $\alpha=0.2$ dB/km) for all distances except 25 km, 50 km, 75 km, 100 km, and 125 km, which employed SMF-28 ULL fiber ($\alpha=0.16$ dB/km). A variable optical attenuator (VOA) is used to adjust the source intensity such that the average intensity at $D_{0}$ is the same for all loop lengths (roughly -25 dBm). We observe a significant correlation between the acoustic level in the laboratory and the magnitude of the intensity fluctuations. Therefore, effort was taken to minimize sounds in the lab while data was collected. Furthermore, to mitigate the effects of vibrations and thermal fluctuations in the environment, all fiber spools are stored inside a polyethylene box placed atop a pneumatically-isolated optical table. The intensity noise PSD was also recorded for the various loop lengths by replacing the oscilloscope with a RFSA. Details on these measurements are found in Appendix B.
@@ V Interference Visibility
Figure 3: (a) Visibility (average of detectors $D_{0}$ and $D_{1}$) as a function of Sagnac loop length for different signal patterns. Curves serve as visual guides. (b) Raw counts registered over 30 s using single-photon detectors when burst-patterning is employed in 200 km fiber ($\phi=0$). The backscattering noise, before it decays, is much stronger than the destructive interference signal ($D_{1}$). Inset shows intensity fluctuations due to fiber phase noise and Poisson noise. Dashed curve is the simulated backscattering noise (peak normalized to experimental data). (c) Counts within pulse detection windows.
We measure interference visibility using the setup in Fig. 1. CW light passes through a 1x1 optical switch driven by a square wave RF signal that determines when it is open or closed, thereby generating a burst pattern. An intensity modulator is used to generate pulses at a repetition rate of 10 MHz with 900 ps pulse width. Inside the Sagnac loop, a phase modulator with a polarizer is used to apply a $\phi=0$ or $\phi=\pi$ phase shift on only one of the interfering paths. A high-speed arbitrary waveform generator (Keysight M8195A) provides the synchronized RF signals to each of the modulators.
@@ VI Discussion
Our work demonstrates that in order to achieve high visibility in long-fiber Sagnac interferometers, a burst-patterning technique may be employed to accommodate backscattering noise. In Sagnac TFQKD, the fiber loop also contains an intensity modulator and a phase modulator for each QKD user [24, 25, 26]. One may show that if there are two users separated by a distance of $L/2$, we may select a burst on-time of $L/(2v_{g})$ and an off-time of $L/v_{g}$ to prevent the backscattering from causing false detections. This burst pattern ensures that the intersection of clockwise and counterclockwise pulses takes place only in the region of the loop where the light from each direction has been attenuated by the loss of a user station. As a result, the intensity difference upon their crossing is low enough that the backscattered light which reaches the detectors at the same time as a signal photon is unlikely to generate a false count. With this configuration, the optimal burst duty cycle is given by [38]
@@
The method of burst-patterning effectively reduces the efficiency of signal generation, in turn reducing the secure key rate. However, all other demonstrated TFQKD systems rely on measuring phase noise, which requires reserving a portion of the communication period for sending reference pulses that do not contribute to key generation. For example, the effective duty cycle for a 200 km channel in Ref. [22] is $25\%$. Therefore, in terms of signal generation efficiency, Sagnac TFQKD is not disadvantaged compared to other configurations.
Our findings help us to assess the fiber length limit of Sagnac TFQKD. Our measurement shows that we can achieve a phase-noise-limited interference visibility of $97\%$ in a 200 km loop, without active phase or polarization compensation. Based on Eq. 3, this visibility corresponds to a phase noise variance $\sigma_{\delta\phi}^{2}=0.06$ rad2, which agrees, within uncertainty, with the value extrapolated from the fit on our measurements in Fig. 2. Ref. [39] shows that in TFQKD we may evaluate the quantum bit error rate, $e$, by the relation $e=\sigma_{\delta\phi}^{2}/4$. Thus, our 200 km interference experiment would yield $e=2\%$, which is sufficient for secure key generation. We note that a 200 km fiber ring would support a communication distance of up to 100 km. Previously, we have also demonstrated a Sagnac TFQKD network with 58 dB overall channel loss simulated with VOAs [26]. Such a system is therefore imminently suitable for settings such as metropolitan quantum networks which consist of long fibers as well as several lossy nodes.
@@ VII Conclusion
In summary, we have characterized the visibility of long-fiber Sagnac interferometers up to a record 200 km in the presence of both phase and backscattering noise, and without active phase or polarization stabilization. We have shown that the backscattering noise problem can be solved with a burst-patterning technique to achieve a phase-noise-limited visibility of $>97\%$ in 200 km fiber. This work can also be extended to assess the scalability of several applications based on long-fiber Sagnac interferometers such as distributed sensing [40], quantum fingerprinting [41], and quantum secret sharing [42].
@@ Appendix A Noise simulation and measurement of backscattering coefficient
Figure 4: (a) Setup used to measure the backscattering coefficient, $\eta$. VOA: variable optical attenuator; C: circulator; SPAD: single-photon avalanche detector; TCSPC: time-correlated single photon counting system. (b) Backscattering response of 20 km SMF-28 fiber for an average input power of -72 dBm (6.5$\times$10-11 W). Data is corrected for the dark count, dead-time, and detection efficiency of the SPAD. Error bars correspond to photon number uncertainty and dark count fluctuation. Black curve is a least-squares fit to Eq. 6.
For reasonable pulse energies, Rayleigh backscattering in fiber can be modelled as a linear time-invariant system. Provided our pulses are sufficiently short compared to the time between the pulses, Eq. 6 gives the system’s impulse response. Therefore, to determine the backscattering from an arbitrary optical signal (e.g., a pulse train), we perform a convolution between Eq. 6 and the signal [44]. In order to account for a loop structure in the fiber, we consider the backscattering response of the clockwise and counterclockwise light independently. Note that the polarization of scattered light is random with respect to the incident light polarization. Therefore, the backscattering noise will on average be evenly distributed between the two detectors in a Sagnac interferometer. In the case that there are inserted components along the fiber, the impulse response of Eq. 6 is modified to include discrete loss points at the appropriate locations.

### v3

# Study of long-fiber Sagnac interferometers for twin field quantum key distribution networks

@@ Abstract
A Sagnac loop structure can help overcome the major difficulty in the practical implementation of a twin field quantum key distribution (TFQKD) network, namely, the need to stabilize the phase of a quantum state over many kilometers of fiber. Unfortunately, Rayleigh backscattering noise limits the signal-to-noise ratio for Sagnac systems containing long fibers and lossy photonic devices. Here, we solve this problem by sending optical pulses in long on-off bursts and using time post-selection on measurements taken with free-run single-photon avalanche detectors. We also investigate the impact of the residual phase noise uncompensated by the Sagnac structure and find that the variance of the phase noise scales as loop length to the third power, verifying an existing calculation in the literature. We measure the interference visibility in Sagnac loops of varying length without active phase stabilization and achieve $>97\%$ visibility in 200 km ultra-low-loss fiber, which is, to our knowledge, the longest fiber Sagnac interferometer demonstrated. Our results suggest that a Sagnac system is feasible for long-distance TFQKD networks, an important step towards the practical implementation of metropolitan quantum networks.
@@ I Introduction
In TFQKD, two remote users (conventionally called Alice and Bob) send coherent pulses encoded with an optical phase to an untrusted central node (conventionally called Charlie) who performs a single-photon interference measurement on the states [16], which requires phase stability between the two interfering optical paths. In most TFQKD experiments, Alice’s and Bob’s laser sources are phase-locked to a reference laser over additional fiber channels, forming a large Mach-Zehnder interferometer [17, 18, 19, 20, 21]. Recently, there have also been demonstrations without phase locking the two sources [22, 23, 24]. For example, Ref. [22] achieves this with the use of strong reference pulses and a fast Fourier transform algorithm to reconcile the phase difference in post-processing. In Ref. [24], a local optical frequency standard is employed at each source to establish an absolute reference, a technique also used in post-measurement pairing QKD [25]. In all of the aforementioned works, however, expensive superconducting nanowire single-photon detectors with precise timing resolution are used in order to measure [17, 19, 21, 22, 24] and/or compensate [18, 20, 23] phase noise. Moreover, it is practically challenging to stabilize interferometers with highly asymmetric long-fiber links, a situation that is inevitable in a multi-user-pair TFQKD network, and this task has only been demonstrated using sub-Hz linewidth lasers with matched frequencies [23].
Unlike the other systems, TFQKD based on a Sagnac interferometer shows promise for easy networking and low-cost implementation [26, 27, 28] due to common-path stable interference. In this configuration, a single laser source (possessed by Charlie) is used to distribute coherent pulses into a fiber ring where one user encodes the clockwise pulses and the other user encodes the counterclockwise pulses. However, Sagnac TFQKD with auto phase stabilization has been restricted to fiber distances of less than 20 km. There are two main challenges to implementing Sagnac TFQKD over long fibers. One is the uncompensated fiber phase noise due to temperature fluctuations and vibrations in the surrounding environment that cause the phase in the fiber to fluctuate during the loop transit time. The other is the Rayleigh backscattering noise in fiber, a challenge also encountered in other TFQKD systems [29, 20] but which is especially problematic in a Sagnac loop due to the bidirectional pulse transmission. This noise is proportional to the intensity and repetition rate of the source at Charlie and thus, for a high-speed and long-distance Sagnac TFQKD network (which contains several lossy photonic devices), the SNR at the detectors becomes limited by the backscattering.
@@
Indeed, no existing literature assesses the feasibility of implementing high-speed long-distance Sagnac TFQKD, with respect to auto phase stability as well as Rayleigh backscattering noise. We overcome both the phase noise and the Rayleigh backscattering noise in fiber to achieve stable interference in loop lengths reaching 200 km. We mitigate the fiber phase noise via thermal and vibrational isolation of the fiber spools; no active phase stabilization is performed, nor do we perform phase post-compensation. Polarization alignment is maintained without fast or automated control. We report measurements on the phase noise as a function of Sagnac loop length, $L$, and find the variance to scale roughly as $L^{3}$. We circumvent the fiber backscattering noise by sending pulses in long, on-off bursts and using post-selection on time-resolved measurements taken with free-run single-photon avalanche detectors (SPADs). We achieve $>99\%$ visibility in 150 km fiber, and $>97\%$ visibility in 200 km. Our results represent record performance in terms of Sagnac interference over long fibers, and suggest that the Sagnac loop is feasible as a platform for long-distance TFQKD networks.
@@ II Sagnac Interferometry in Fiber
Figure 1: Sagnac interferometer experimental setup. Light from a continuous-wave laser launches into a fiber ring of length $L$ via a 50:50 beam splitter (BS). Interference of the counterpropagating paths is measured by detectors $D_{0}$ and $D_{1}$. Components on the hexagonal or square platforms are added when measuring phase noise or visibility, respectively. Classical photodetectors and single-photon avalanche detectors are used for phase noise and visibility measurements, respectively. IM: intensity modulator; VOA: variable optical attenuator; C: circulator; PC: polarization controller; PM: phase modulator; RFSA: rf spectrum analyzer; Osc: oscilloscope.
@@ IV Fiber Phase Noise
$I(\delta\phi(t))$ was measured for loop lengths ranging from five to 125 km. Spools consisted of SMF-28 fiber (fiber attenuation coefficient, $\alpha=0.2$ dB/km) for all distances except 25 km, 50 km, 75 km, 100 km, and 125 km, which employed SMF-28 ULL fiber ($\alpha=0.16$ dB/km). A variable optical attenuator (VOA) is used to adjust the source intensity such that the average intensity at $D_{0}$ is the same for all loop lengths (roughly -25 dBm). We observe a significant correlation between the acoustic level in the laboratory and the magnitude of the intensity fluctuations. Therefore, effort was taken to minimize sounds in the lab while data was collected. Furthermore, to mitigate the effects of vibrations and thermal fluctuations in the environment, all fiber spools are stored inside a polyethylene box placed atop a pneumatically-isolated optical table. In a real-world Sagnac TFQKD network, fibers would likely be deployed underground, which may similarly assist in mitigating these noise effects as was observed in Ref. [32]. The intensity noise PSD was also recorded for the various loop lengths by replacing the oscilloscope with a RFSA. Details on these measurements are found in Appendix B.
@@ V Interference Visibility
Figure 3: (a) Visibility (average of detectors $D_{0}$ and $D_{1}$) as a function of Sagnac loop length for different signal patterns. Curves serve as visual guides. (b) Counts registered over 30 s using single-photon detectors when burst-patterning is employed in 200 km fiber ($\phi=0$). A moving average over a 1 $\mu$s window is used. The backscattering noise, before it decays, is much stronger than the destructive interference signal ($D_{1}$). Inset shows intensity fluctuations due to fiber phase noise and Poisson noise. Dashed curve is the simulated backscattering noise (peak normalized to experimental data). (c) Raw counts within pulse detection windows.
We measure interference visibility using the setup in Fig. 1. CW light passes through a 1x1 optical switch (extinction ratio $>25$ dB) driven by a square wave RF signal that determines when it is open or closed, thereby generating a burst pattern. An intensity modulator is used to generate pulses at a repetition rate of 10 MHz with 900 ps pulse width. Inside the Sagnac loop, a phase modulator with a polarizer is used to apply a $\phi=0$ or $\phi=\pi$ phase shift on only one of the interfering paths. A high-speed arbitrary waveform generator (Keysight M8195A) provides the synchronized RF signals to each of the modulators.
@@ VI Discussion
Our work demonstrates that in order to achieve high visibility in long-fiber Sagnac interferometers, a burst-patterning technique may be employed to accommodate backscattering noise. In Sagnac TFQKD, the fiber loop also contains an intensity modulator and a phase modulator for each QKD user [26, 27, 28]. One may show that if there are two users separated by a distance of $L/2$, we may select a burst on-time of $L/(2v_{g})$ and an off-time of $L/v_{g}$, where $v_{g}$ is the group velocity of light in fiber, to prevent the backscattering from causing false detections. This burst pattern ensures that the intersection of clockwise and counterclockwise pulses takes place only in the region of the loop where the light from each direction has been attenuated by the loss of a user station. As a result, the intensity difference upon their crossing is low enough that the backscattered light which reaches the detectors at the same time as a signal photon is unlikely to generate a false count. With this configuration, the optimal burst duty cycle is given by [41]
@@
The method of burst-patterning effectively reduces the efficiency of signal generation, in turn reducing the secure key rate. However, all other demonstrated TFQKD systems rely on measuring phase noise, which requires reserving a portion of the communication period for sending reference pulses that do not contribute to key generation. For example, the effective duty cycle for a 200 km channel in Ref. [22] is $25\%$. Therefore, in terms of signal generation efficiency, Sagnac TFQKD is not disadvantaged compared to other configurations. Furthermore, Rayleigh scattering limits the SNR even in one-way TFQKD systems [29, 20], albeit at much longer transmission distances. This effect is due to double Rayleigh scattering of the strong reference pulses used for phase estimation. In such cases, the use of wavelength division multiplexing to allow for setting the wavelength of the reference light different from the quantum signal is needed to significantly reduce the double Rayleigh scattering noise [18, 21].
Our findings help us to assess the fiber length limit of Sagnac TFQKD. Our measurement shows that we can achieve a phase-noise-limited interference visibility of $97\%$ in a 200 km loop, without active phase or polarization compensation. Based on Eq. 3, this visibility corresponds to a phase noise variance $\sigma_{\delta\phi}^{2}=0.06$ rad2, which agrees, within uncertainty, with the value extrapolated from the fit on our measurements in Fig. 2. Ref. [42] shows that in TFQKD we may evaluate the quantum bit error rate, $e$, by the relation $e=\sigma_{\delta\phi}^{2}/4$. Thus, our 200 km interference experiment would yield $e=2\%$, which is sufficient for secure key generation. We note that a 200 km fiber ring would support a communication distance of up to 100 km. Previously, we have also demonstrated a Sagnac TFQKD network where three users were placed along a fiber loop with 58 dB overall channel loss simulated with VOAs [28] (equivalent to 362 km SMF-28 ULL fiber). An asymptotic key rate of $R=2.014\times 10^{-6}$ was obtained, comparable to the performance of other configurations (e.g., Ref. [18] obtain $R=8.527\times 10^{-7}$ over a 369 km channel). Therefore, while our system with auto phase stability is unlikely to achieve transmission distances as long as other setups for TFQKD, it can achieve similar key rates over equivalent channel losses, with the unique advantages of being low-cost and scalable to many users. This characterization suggests that such a system is feasible for settings such as ring metropolitan quantum networks [43] which consist of long fibers as well as several lossy nodes.
Lastly, we note that in Sagnac TFQKD where users receive and modulate light originating from an untrusted source, additional components should be implemented at each user station to guarantee security [44, 45]. As discussed in Ref. [26], taps, photodetectors, and bandpass filters are necessary for Alice and Bob to detect and limit strong optical injections from the outside, so as to prevent eavesdroppers from probing the sources. A VOA is used to attenuate the pulses traveling back to Charlie to single-photon level. Thus, it is possible to achieve unconditional security with the Sagnac configuration for TFQKD while also maintaining inexpensive hardware requirements.
@@ VII Conclusion
In summary, we have characterized the visibility of long-fiber Sagnac interferometers up to a record 200 km in the presence of both phase and backscattering noise, and without active phase stabilization. We have shown that the backscattering noise problem can be solved with a burst-patterning technique to achieve a phase-noise-limited visibility of $>97\%$ in 200 km fiber. This work can also be extended to assess the scalability of several applications based on long-fiber Sagnac interferometers such as distributed sensing [46], quantum fingerprinting [47], and quantum secret sharing [48].
@@ Appendix A Noise simulation and measurement of backscattering coefficient
Figure 4: (a) Setup used to measure the backscattering coefficient, $\eta$. VOA: variable optical attenuator; C: circulator; SPAD: single-photon avalanche detector; TCSPC: time-correlated single photon counting system. (b) Backscattering response of 20 km SMF-28 fiber for an average input power of -72 dBm (6.5$\times$10-11 W). Data is corrected for the dark count, dead-time, and detection efficiency of the SPAD. Error bars correspond to photon number uncertainty and dark count fluctuation and are plotted on a fraction of data points for visual clarity. Black curve is a least-squares fit to Eq. 6.
For reasonable pulse energies, Rayleigh backscattering in fiber can be modeled as a linear time-invariant system. Provided our pulses are sufficiently short compared to the time between the pulses, Eq. 6 gives the system’s impulse response. Therefore, to determine the backscattering from an arbitrary optical signal (e.g., a pulse train), we perform a convolution between Eq. 6 and the signal [50]. In order to account for a loop structure in the fiber, we consider the backscattering response of the clockwise and counterclockwise light independently. Note that the polarization of scattered light is random with respect to the incident light polarization. Therefore, the backscattering noise will on average be evenly distributed between the two detectors in a Sagnac interferometer. In the case that there are inserted components along the fiber, the impulse response of Eq. 6 is modified to include discrete loss points at the appropriate locations.